OpenStudy (anonymous):

Find the zeros of the function: f(x) = x2 - 7x. Enter the largest zero value.

OpenStudy (bahrom7893):

factor out an x and set the equation = 0

OpenStudy (bahrom7893):

x(x-7)=0 x1 = 0; x2 - 7 = 0 x1 = 0; x2 = 7

OpenStudy (bahrom7893):

which one's bigger?

OpenStudy (nenadmatematika):

x*(x-7)=0 x=0 or x=7 so answer is 7

jhonyy9 (jhonyy9):

x^2 -7x=0 x(x-7)=0 x_1=0 x-7=0 x_2=7 x_1=0 x_2=7 these are roots of

OpenStudy (anonymous):

thank you bahrom7893! (: